A mass of 1 kg stretches a spring 39.2 cm. The means by which the mass moves offers a damping force equal to 8 times the instantaneous speed. Find the equation of motion if the mass is released from equilibrium with a descending velocity of 5 m / s.
Answer #1

Given mas ma Ikg L- 39.2 cm 39.2410-2m c=8 9-9.8 m/s الها mg 1x9.8 ka = 25 L L 39.2x10-2 *100=0, x(o)= smis equation is matsubstitute (i=0 in @ -) (2= 25 5=32 ) in substitute ci=0, ₂ = e-4t sinst ut) = win